OpenStudy (anonymous):

5x+10y=0 & -4x+10y=-36 written in matrix format -> {{5,10,0}, {-4,10,-36}} and I need to parameterize it using x and y, so in the format x{{#},{#}} + y{{#},{#}} = {{0},{36}}.

OpenStudy (arabpride):

This isn't linear algebra..? cx

OpenStudy (doc.brown):

\[\left[\begin{matrix}5 & 10 & 0 \\ -4 & 10 & -36\end{matrix}\right]\]\[\left[\begin{matrix}1 & 2 & 0 \\ -4 & 10 & -36\end{matrix}\right]\]\[\left[\begin{matrix}1 & 2 & 0 \\ 0 & 18 & -36\end{matrix}\right]\]\[\left[\begin{matrix}1 & 2 & 0 \\ 0 & 1 & -2\end{matrix}\right]\]\[\left[\begin{matrix}1 & 0 & -4 \\ 0 & 1 & -2\end{matrix}\right]\]